Week of 170529

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Julia (WLCG), Maarten (Alice,WLCG), Kate (WLCG, DB, chair), Ben (computing), Marian (network), Andrea M (MW)
  • remote: David (IN2P3), Luca (CNAF), Ulf (NDGF), Stefano (CMS), Pepe (PIC), Vincenzo (EGI), Dimitri (KIT), Victor (JINR, excused)

Experiments round table:

  • CMS reports ( raw view) -
    • Load on computing grid still low due to lack of requests
    • No major or new problem this week

  • ALICE -
    • Continued very high activity
      • New record of 143k running jobs this morning

  • LHCb reports ( raw view) -
    • Activity
      • User analysis and Monte Carlo simulation
      • Stripping v24 is almost finished.
    • Site Issues

Sites / Services round table:

  • ASGC: nc
  • BNL: NTR
  • CNAF:
    • IPV6: tests on gpfs cluster ongoing.
    • CMS: ce01-lcg has been dismissed but receives tests, why?
Maarten commented that the CMS VO feed hasn't been updated and a ticket to VO support can be opened. Marian commented the VO feed is taken directly from the pilot factories.
  • EGI: ntr
  • FNAL: NTR
  • IN2P3: ntr
  • JINR:
    • Network glitch at 24-05-2017.
    • Working on IPv6 routing on external links, LAN is ready for IPv6.
  • KISTI: nc
  • KIT: ntr
  • NDGF: one ALICE pool has been down for two days, no complaints from ALICE
  • NL-T1: nc
  • NRC-KI: nc
  • OSG: nc
  • PIC: 20% of worker nodes have been moved to HTCondor. Remaining capacity for LHC experiments will be moved in the coming weeks
  • RAL: nc
  • TRIUMF: NTR

  • CERN computing services: ntr
  • CERN storage services: nc
  • CERN databases:
    • CSMONR ADG database not up to date (OTG:0037748). Issues related to a bug in ADG, SR with Oracle has been opened
    • CMSONR firewall blocking new DB connections for 90mins due to misconfiguration on Friday (OTG:0037788)
  • GGUS:
    • Wed this week there will be a regular GGUS release with the usual alarm tests
    • NOTE: for the installation of OS patches and an upgrade of the Remedy back-end
      there will be a downtime 06:00-10:00 UTC
  • Monitoring: nc
  • MW Officer: NTR
  • Networks: NTR
  • Security: NTR

AOB:

  • ATTENTION: next meeting on Tuesday June 6 !
Edit | Attach | Watch | Print version | History: r15 < r14 < r13 < r12 < r11 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r15 - 2017-05-29 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback